OpenStudy (anonymous):

(2a 2/3) raised to 2 is the correct answer 4a4/3 ???????????

OpenStudy (anonymous):

the denominator should be 9, 3^2=9

OpenStudy (anonymous):

recall : (a/b)^n=a^n/b^n

OpenStudy (anonymous):

sooo 4a4/9 ??????

OpenStudy (anonymous):

you got it

OpenStudy (anonymous):

sure??? lol

OpenStudy (anonymous):

no man, i am just guessing

OpenStudy (anonymous):

your answer was wrong you mutiply 2*2/3 and get 4/3

OpenStudy (anonymous):

oh ****, my bad

OpenStudy (anonymous):

well, you learn from your failures

OpenStudy (anonymous):

its fine....thats why i questioned it

OpenStudy (anonymous):

but th

OpenStudy (anonymous):

thanks